A uniform, solid sphere of radius 5 cm and mass 4.75 kg starts with a purely translational speed of 3.75 m/s at the top of an inclined plane. The surface of the incline is 1m long, and is tilted at an angle of 22 degrees with respect to the horizontal. Assuming the sphere rolls without slipping down the incline, calculate the sphere's final translational speed  at the bottom of the ramp.
Answer #1

using conservation of energy

pE at top + translational kE at top = translational kE at bottom + rotational kE at bottom

mg (d sin x) + 0.5 m u^2 = 0.5 m v^2 + 0.5* I w^2

mg d sin x + 0.5"m u^2 = 0.5 m v^2 + 0.5* (2/5) m r^2* ( v^2/r^2)

9.8* 1* sin 22 + 0.5* 3.75^2 = 0.5* v^2 + 0.5* (2/5)* v^2

9.8 sin 22 + 0.5* 3.75^2 = 0.7 v^2

v = 3.91 m/s

======

Comment before rate in case any doubt, will reply for sure.. goodluck